# On the Utility of Domain-Adjacent Fine-Tuned Model Ensembles for Few-shot Problems

### v1

@@ 3.2.2 Comparison of FT and DA(FT)2
Dataset 2-shot 4-shot 8-shot 16-shot 32-shot 64-shot 128-shot
MRPC 45.30 43.69 46.97 3.26 32.97 25.40 24.17
QQP 78.40 79.36 94.01 92.48 71.76 52.46 17.84
STSB 49.33 51.92 53.74 57.91 51.06 32.28 3.05
Here we explore the few-shot case when some training data is available from the target domain. The generalized approach with any training data is to fine-tune a model with the data so that the model can perform the domain specific task. The key research question is: should we fine-tune a base model (i.e., FT in Figure 1), or fine-tune a DAFT (i.e., DA(FT)2 in Figure 1) model?  [^4] In our fine-tuning experiments, we found that only updating the weights of the header layer does not guarantee good performance and was easily surpassed by different models designed to perform similar tasks without any fine-tuning (zero-shot). Hence, we chose to do all our performance analysis by focusing on fully fine-tuned models. The performance comparison of updating the weights of the entire model versus only the header layer is provided in the Appendix (Table 6).
Table 1 provides the average performance comparison of the FT and DA(FT)2 models using the Relative Improvement ($RI$) metric when the number of samples to fine-tune FT and DA(FT)2 is varied from $n=2$ to $256$ (The full results for both tasks are in Appendix, Table 6 and 7). The values in the table are the average of $RI$ of DA(FT)2 compared to FT the Roberta base model, when all the DAFT models also have Roberta as the base model. A positive (negative) value of $RI$ means DA(FT)2 is performing better (worse) compared to FT, and the larger (smaller) the value the better (worse) DA(FT)2 is performing. We observe that DA(FT)2 outperforms FT by a large margin when $n$ is within the range of $2$ to $64$. We see only one negative value in the table when $n=256$, which indicates that FT is outperforming DA(FT)2 in most cases. The presence of almost all positive values of $RI$ showcases the importance of using a fine-tuned DAFT instead of FT when the sample size is small.
@@
[linecolor=black!0,backgroundcolor=orange!10]These results indicate that few-shot fine-tuning of DAFT models is generally more beneficial than fine-tuning base models.
The performance of DAFT models showcases the significance of using them for domain adjacent tasks. From a practical standpoint, a major advantage of using DAFT models is that we get them for free. However, as our results show, there is considerable uncertainty about what is the right DAFT model to choose (especially in the zero-shot setting), as the performance can significantly vary between different DAFT models. In the next section, we explore the use of ensemble methods to address this model selection problem.
@@ 4 Ensembling of DAFT Models
In this section, we evaluate the performance of the DAFT models using two distinct Ensembling methods and benchmark their performances.
4.1 DAFT-E ${}^{\text{Z}}$ Performance
Figure 4: RI of DAFT ${}^{\text{Z}}$ over DAFT-E ${}^{\text{Z}}$: For each test dataset on the sentiment analysis task, we consider 15 DAFT LLMs (fine-tuned on data different from the test data), and consider the RI of DAFT ${}^{\text{Z}}$ over the average ensemble DAFT-E ${}^{\text{Z}}$ of the 15 DAFT LLMs. Values less than 0 indicate performance degradation. The red squares $\blacksquare$ denote the performance of zero-shot prompting on larger LLMs (Bart-large-mnli, Roberta-large-mnli, and Opt-1.3b). The results show that DAFT-E ${}^{\text{Z}}$ usually has significantly better zero-shot performance than DAFT ${}^{\text{Z}}$. However, there are some cases where some specific DAFT models can outperform the average ensemble (leading to $\text{RI}>0$).
@@ 4.1 DAFT-E ${}^{\text{Z}}$ Performance
The average Ensemble (DAFT-E ${}^{\text{Z}}$) method is a zero-shot method that can be used as an alternative to randomly picking one of the DAFT models. In DAFT-E ${}^{\text{Z}}$, the output probabilities from all the DAFT models are averaged and the decision is taken on that value (max-voting is another feasible approach). To perform average ensembling, we run inference on all the DAFT models. However, since these DAFT models do not need any further training or fine-tuning and inference can be run in parallel, it is feasible to utilize them efficiently with the abundance of computational devices (mostly low end) available today. Figure 4 shows the Relative Improvement of all the 15 DAFT models and the three larger models (as previously discussed in 3.1.2) for sentiment analysis, compared to the average ensemble of the DAFT models, i.e., DAFT-E ${}^{\text{Z}}$.
@@
From Figure 4 we see that very few points have a positive $RI$ value , implying that the average ensemble method (DAFT-E ${}^{\text{Z}}$) is better than most of the DAFT models. Moreover, the performance of the larger LLMs (red squares) are not as good as DAFT-E ${}^{\text{Z}}$. Furthermore, for the Tweet dataset, for which none of the DAFT models performed well and was beaten in performance by two of the larger models (Bart-large-mnli, Roberta-large-mnli), we now have $RI$ values close to zero. Hence, DAFT-E ${}^{\text{Z}}$ has a performance very close to those larger models and shows the importance of using an ensemble even when individual DAFT models cannot perform well. For the text similarity task we see a similar trend in the performance of DAFT-E ${}^{\text{Z}}$ and the $RI$ figure is in the Appendix (Fig: 15). In Section 5, we argue theoretically (Proposition 1) that it is always better to use the average ensemble method over choosing a DAFT model randomly from the set of DAFTs ($\mathcal{M}$).
{mdframed}
@@
[linecolor=black!0,backgroundcolor=orange!10]For zero-shot tasks, the ensemble of the DAFT models (DAFT-E ${}^{\text{Z}}$) is on average a better choice than any individual DAFT model.
@@
The average ensemble is therefore a great data-agnostic solution; however, when training data from the target domain is available, we show the data-informed ensembling can be a stronger choice. We thus pursue a weighted ensemble of the DAFT models (DAFT-E) in the following section.
4.2 Comparison of DAFT-E and DA(FT)2
@@
Table 2: Average $RI$ of DA(FT)2 compared to DAFT-E (Top: 6 sentiment analysis tasks. Bottom: 3 text similarity tasks). Negative values indicate that DAFT-E is better than DA(FT)2; larger values imply higher improvements. Table 8 and 9 in the Appendix gives more details. In the very few-shot setting, DAFT-E outperforms DA(FT)2 in almost all cases.
Dataset 2-shot 4-shot 8-shot 16-shot 32-shot 64-shot 128-shot
@@
Dataset 2-shot 4-shot 8-shot 16-shot 32-shot 64-shot 128-shot
MRPC -4.91 -3.82 -4.83 -4.96 -2.52 3.12 7.57
QQP 0.22 -0.31 4.70 2.89 5.96 5.14 8.07
STSB 11.30 5.39 22.66 6.18 1.09 2.56 4.78
@@
We define DAFT-E as the weighted ensemble method, where the weights of the ensemble layer are learned using the training data from the target dataset. To perform the weighted ensemble in DAFT-E we utilized two specific regression methods, a) Random forest-based regression ($RF$) and b) SGD-based linear regression ($LR$). The parameters of these methods are discussed in A.5 of the Appendix. We observe the following: (i) $LR$ always has better performance and less variation in performance for smaller sample data, (ii) for higher sample data, i.e., $n$>32, $RF$ usually performs very similar to $LR$, (iii) for sentiment analysis tasks, $LR$ and $RF$ both perform well with similar performance at higher sample data, and (iv for textual similarity tasks, $LR$ tends to show minimal improvement in performance with more samples, whereas $RF$ shows promising results. Based on these observations, and $LR$ being the most lightweight, we show the performance of DAFT-E with $LR$ here and discuss more on the results in the Appendix (Table 8 and 9) and in Section 7 (Limitations).
@@
In the previous section, we showed that DA(FT)2 usually performs better than FT for few-shot learning. Hence, we compare the performance of DAFT-E (weighted ensemble) with DA(FT)2. For a fair comparison, the same amount of data is used when fine-tuning DA(FT)2 and learning the weights of the ensemble layer. Table 2 shows the average RI of DA(FT)2 compared to DAFT-E. Hence, a $RI$ of positive (negative) means the DA(FT)2 performed better (worse, resp.) compared to DAFT-E. From Table 2 it is evident that DAFT-E outperforms DA(FT)2 for sentiment analysis tasks, i.e., all values are negative in the table. On the other hand, for textual similarity DAFT-E usually performs well for smaller data samples (except STSB, see Appendix A.3.2 for discussion), but underperforms on more-data (positive $RI$ values). However, $RF$ performed well with more samples and actually performs better or similarly to DA(FT)2 for $n>16$ (Table 10 in the Appendix). This shows that weighted ensembling is effective, with task-specific variations between $RF$ and $LR$.
Note that updating the ensemble layer depends on the predictions of the DAFT models, and requires solving either (i) a simple linear equation to optimize the weights of the ensemble layer or (ii) running a lightweight $RF$ (max depth = $2$). Thus, only a single inference on each of the DAFT models with the few-shot samples is necessary. In contrast, for fine-tuning a DAFT model, we need to perform back-propagation through the large model to update weights, which is computationally expensive.
@@
[linecolor=black!0,backgroundcolor=orange!10]For few shot tasks, weighted ensembling of DAFT models (DAFT-E) can outperform DA(FT)2, with a suitable ensembling method.
@@ 4.3 Overall Comparison
(a) Amazon
(b) Cornell
(c) IMDB
(d) SST2
(e) Tweet
(f) Yelp
Figure 5: Performance comparison of DAFT-E ${}^{\text{Z}}$, DAFT-E, single-best DAFT and FFT. The error interval for DAFT-E is based on the random choice of the few-shot samples used to learn the weights of the ensemble aggregated over 10 trials. The single best DAFT for different datasets are: Amazon (Roberta - Yelp), Cornell (Roberta - SST2), IMDB (xlnet - amazon), SST2 (xlnet - Cornell), Tweet (xlnet - SST2), Yelp (xlnet - amazon).
For benchmarking purposes, let us assume that we have the entire training data ($D_{T}$) of the target domain, and have a model that is fine-tuned on $D_{T}$ (FFT on $D_{T}$). Let us compare the performance of the ensemble methods with this FFT on different datasets. Figure 5(a)-5(f) show the performance comparison of DAFT-E ${}^{\text{Z}}$ (green dash-dot line), DAFT-E (purple dash-dot line), (single) best DAFT (black dotted line), and FFT on $D_{T}$ (red dotted line) for all the six datasets of the sentiment analysis task. The result for text similarity tasks is provided in the Appendix (Figure 16). From Figure 5, we observe the advantage of using DAFT-E over DAFT-E ${}^{\text{Z}}$ when training data is available. It is interesting to note that DAFT-E has a strong increasing trend in performance (against fine-tuning data size) in most cases, and with the increase of fine-tuning data catches up with either the single best DAFT or FFT for the target domain. The performance of DAFT-E compared to the best solution, i.e., FFT on the target domain, can be bounded theoretically and is given by Proposition 2 in Section 5. Table 3 shows the overall comparison of the computational cost of all the five options that we have discussed in this paper. It should be noted that none of the ensemble methods suffer from the high computational complexity of back-propagation computation; further, using sparse weighting we can reduce the inference cost in DAFT-E compared to DAFT-E ${}^{\text{Z}}$. {mdframed}[linecolor=black!0,backgroundcolor=orange!10]DAFT-E ${}^{\text{Z}}$ and DAFT-E are strong lightweight methods for leveraging already available DAFT models, often matching the single best DAFT model (which is only known post-hoc), and at times matching the best possible performance of full fine-tuning with a large fine-tuning set.
Additional results.
Lastly, DAFT-E ${}^{\text{Z}}$ has shown favorable performance compared to a state-of-the-art zero shot ensemble model, Uniform Soup (Model Soup) Wortsman et al. (2022), and the calculation of LEEP scores Nguyen et al. (2020) on transferability also matches the performance heat map showed in Figure 2. Details of these results with discussion are in Appendix (Tables 11, 12 and Figure 17). Moreover, a table comparing the fundamental concepts of DAFT-E and DAFT ${}^{\text{Z}}$ to the model soup is provided in Table 13
Table 3: Computational cost of methods shown in Figure 1. $n$ denotes the number of few-shot samples. $C_{F}$ and $C_{B}$ denote the computational costs of a forward and backward pass for a LLM. $E$ is the number of fine-tuning epochs. $N$ is the number of DAFT models available to ensemble, $\bar{N}\leq N$ is the number of nonzero weights in the weighted ensemble.
@@
DAFT ${}^{\text{Z}}$ 00 $C_{F}$
DAFT-E ${}^{\text{Z}}$ 00 $N\cdot C_{F}$
@@
DAFT-E $n(C_{F}+E)$ $\bar{N}\cdot C_{F}$

### v2

@@ 3.2.2 Few Shot Comparison of FT & DA(FT)2
MRPC 44.88 42.92 48.08 46.74 32.15 23.47 20.90
QQP 72.93 75.17 84.44 86.97 65.36 48.50 13.86
STSB 48.36 49.31 49.52 53.08 49.88 29.24 -0.12
When (some) training data is available from the target domain, the general approach is to fine-tune a model with that data to perform the domain specific task. The key research question is: should we fine-tune a base model (FT in Fig. 1), or fine-tune a DAFT (DA(FT)2 in Fig. 1)? We explored two scenarios, with weight updates of i) only the header layer connection and ii) the whole model, and decided to go with the weight update of the whole model 555 We found that updating the weights of the header layer only does not guarantee good performance and can be surpassed by models designed to perform similar tasks without any fine-tuning (zero-shot). So, we chose to do all performance analysis with fully fine-tuned models. Performance comparison of updating the weights of the entire model versus only the header layer is provided in the Appendix (Table 9).. Table 2 provides the average performance comparison of the FT and DA(FT)2 models using the Relative Improvement ($RI$) metric when the number of samples to fine-tune is varied from $n=2$ to $128$ (Full results are in Appendix, Table 9 and 10)  [^6] The $RI$ values in the table are generated with Roberta as the base models for both DA(FT)2 and FT.. A positive (negative) value of $RI$ means DA(FT)2 is performing better (worse) compared to FT, and the larger (smaller) the value the better (worse) DA(FT)2 is performing. We observe that DA(FT)2 outperforms FT by a large margin when $n$ is small ($2$ to $64$), and see only one negative value in the table ($n=128$). This indicates that DA(FT)2 is outperforming FT in most (few shot) cases on the target task. Please note that if there are a moderate amount of data-samples (for the current tasks, $n>128$) available from the target domain (not a data-scarce environment) FT is expected to perform similarly as DA(FT)2.
@@
[linecolor=black!0,backgroundcolor=orange!10]For few-shot tasks, fine-tuning DAFT models, that is DA(FT)2, is generally more beneficial than fine-tuning base models (FT).
Figure 5: RI of DAFTs over DAFT-E ${}^{\text{\!\scriptsize Z}}$ for sentiment analysis: For each test dataset, we consider 15 DAFT LLMs (fine-tuned on data different from the test data), and consider the RI of DAFT ${}^{\text{\!\scriptsize Z}}$ over DAFT-E ${}^{\text{\!\scriptsize Z}}$ of 15 DAFT LLMs. Values less than 0 indicate performance degradation. The red squares $\blacksquare$ denote the performance of zero-shot classification on larger LLMs (Bart-large-mnli, Roberta-large-mnli), and prompting with Opt-1.3b. The results show that DAFT-E ${}^{\text{\!\scriptsize Z}}$ usually has significantly better zero-shot performance than DAFT ${}^{\text{\!\scriptsize Z}}$. However, there are some cases where some specific DAFT models can outperform the average ensemble (leading to $\text{RI}>0$).
From a practical standpoint, a major advantage of using DAFT models is that we get them for free. However, as our results show, there is considerable uncertainty about what is the right DAFT model to choose (especially in the zero-shot setting), as the performance can significantly vary between different DAFT models (Fig. 3). Also, note that while DAFT ${}^{\text{\!\scriptsize Z}}$ is zero shot, DA(FT)2 is computationally expensive, needing back-propagation. Hence, in the next section, we explore the use of ensemble methods, which is a cheaper alternative and addresses the model selection problem.
@@ 4 Ensembling of DAFT Models
In this section, we evaluate the performance of the DAFT models using two distinct Ensemble methods and benchmark their performances. Advantages of using ensemble methods are discussed in Appendix A.5.
4.1 DAFT-E ${}^{\text{\!\scriptsize Z}}$ Performance (Zero Shot)
@@
The average Ensemble (DAFT-E ${}^{\text{\!\scriptsize Z}}$) method is a zero-shot method that can be used as an alternative to randomly picking one of the DAFT models. In DAFT-E ${}^{\text{\!\scriptsize Z}}$, the output probabilities from all the DAFT models are averaged and the decision is taken on that value (max-voting is another feasible approach). To perform DAFT-E ${}^{\text{\!\scriptsize Z}}$, we run inference on all the DAFT models. However, since these DAFT models do not need any further training or fine-tuning and inference can be run in parallel, it is feasible to utilize them efficiently with the abundance of computational devices (mostly low end) available today.
@@
Fig. 5 shows the Relative Improvement of all the 15 DAFT models and the three larger models (as previously discussed in 3.1.2) for sentiment analysis, compared to the average ensemble of the DAFT models, i.e., DAFT-E ${}^{\text{\!\scriptsize Z}}$. There are very few cases ($5/90$ for DAFT ${}^{\text{\!\scriptsize Z}}$, and $1/18$ for larger LLMs) in Fig. 5 where a positive $RI$ value is seen, implying that the average ensemble method (DAFT-E ${}^{\text{\!\scriptsize Z}}$) is better than most of the DAFT models. Moreover, the performance of the larger LLMs (red squares) are not as good as DAFT-E ${}^{\text{\!\scriptsize Z}}$. Furthermore, for the Tweet dataset, for which none of the DAFT models performed well and was beaten in performance by two of the larger models (Bart-large-mnli, Roberta-large-mnli), we now have $RI$ values close to zero. Furthermore, for the zero-shot Tweet task, while the larger models outperformed all the DAFT ${}^{\text{\!\scriptsize Z}}$ models (Fig 5), DAFT-E ${}^{\text{\!\scriptsize Z}}$ is able to match their performance, highlighting the utility of ensembles. Hence, DAFT-E ${}^{\text{\!\scriptsize Z}}$ has a performance very close to those larger models and shows the importance of using an ensemble even when individual DAFT models cannot perform well. For text similarity task we see a similar trend in the $RI$ performance (Fig. 6).
Figure 6: Relative Improvement of DAFTs compared to DAFT-E ${}^{\text{\!\scriptsize Z}}$ for Text similarity task: For each test dataset, we consider 6 DAFT FMs (fine-tuned on data different from the test data), and consider the RI of DAFT ${}^{\text{\!\scriptsize Z}}$ over the DAFT-E ${}^{\text{\!\scriptsize Z}}$. Values less than 0 indicate better performance of DAFT-E ${}^{\text{\!\scriptsize Z}}$.
Lastly, to evaluate the performance of DAFT-E ${}^{\text{\!\scriptsize Z}}$ on practical and modern LLM benchmark datasets, we performed an ensemble of 8 DAFT models across the 9 datasets selected for the target task of Medical Knowledge (see Section 3.2 for details on the datasets). The results are shown is Fig. 7. We observe that DAFT-E ${}^{\text{\!\scriptsize Z}}$ outperforms the base model (Llama 3.1) on all instances (red squares with black border) and surpasses individual DAFT models in 56 out of 72 cases (circles in the figure). In Section 5, we argue theoretically (Proposition 1) that it is better (in an expected sense) to use DAFT-E ${}^{\text{\!\scriptsize Z}}$ over choosing a DAFT model randomly from the set of DAFTs.
@@
{mdframed} [ userdefinedwidth= 8.5cm, align=left, linecolor=black!0,backgroundcolor=orange!10]For zero-shot tasks, just a simple ensemble of DAFT models (DAFT-E ${}^{\text{\!\scriptsize Z}}$) is on average a better choice than any individual DAFT model.
@@
The average ensemble is therefore a great data-agnostic solution; however, when training data from the target domain is available, we show that the data-informed ensembling can be a stronger choice. We thus pursue a weighted ensemble of the DAFT models (DAFT-E) in the following section.
4.2 Few-Shot DAFT-E and DA(FT)2
@@
Table 3: Average $RI$ of DA(FT)2 compared to DAFT-E (Top 6: sentiment analysis tasks. Bottom 3: text similarity tasks). Negative values indicate that DAFT-E is better than DA(FT)2 on target (task); larger values imply higher improvements. Table 11 and 12 in the Appendix gives more details. In the very few-shot setting, DAFT-E outperforms DA(FT)2 in almost all cases.
Target 2-shot 4-shot 8-shot 16-shot 32-shot 64-shot 128-shot
@@
MRPC -5.19 -4.34 -4.11 -3.60 -3.12 1.53 4.74
QQP -2.85 -2.64 -0.46 -0.06 2.01 2.41 4.42
STSB -5.29 -5.77 -4.07 -1.89 -1.97 0.57 1.28
@@
We define DAFT-E as the weighted ensemble method, where the weights of the ensemble layer are learned using the training data from the target dataset. To perform the weighted ensemble in DAFT-E we utilized two specific regression methods, a) Random forest-based regression ($RF$) and b) SGD-based linear regression ($LR$) [^7] Parameters of these methods are given in Appendix A.6.. Since, $LR$ is more lightweight with comparable or better performance than $RF$, we considered $LR$ to generate the results of DAFT-E (More discussion on $LR$ and $RF$ performance comparison with the results are in Appendix: Table 11 and 12).
Figure 7: Relative Improvement of DAFTs compared to DAFT-E ${}^{\text{\!\scriptsize Z}}$ for Medical Knowledge task: For each dataset, we consider 8 DAFTs, and consider the RI of DAFT ${}^{\text{\!\scriptsize Z}}$ over the DAFT-E ${}^{\text{\!\scriptsize Z}}$. Values less than 0 indicate better performance of DAFT-E ${}^{\text{\!\scriptsize Z}}$.
@@
In Section 3.2.2, we observed that DA(FT)2 usually performs better than FT for few-shot learning. Hence, we compare the performance of DAFT-E (weighted ensemble) with DA(FT)2. For a fair comparison, the same amount of data is used when fine-tuning DA(FT)2 and learning the weights of the ensemble layer. Table 3 shows the average RI of DA(FT)2 compared to DAFT-E. An $RI$ of positive (negative) means that DA(FT)2 performed better (worse, resp.) compared to DAFT-E. From Table 3 it is evident that DAFT-E outperforms DA(FT)2 for sentiment analysis tasks, i.e., all values are negative in the table. On the other hand, for textual similarity DAFT-E usually performs well for smaller data samples but under-performs on more-data (positive $RI$ values). However, $RF$ performed well with more samples and actually performs better or similarly to DA(FT)2 for $n>16$ (Table 13 in the Appendix). This shows that weighted ensembling is effective, with task-specific variations between $RF$ and $LR$. Note that updating the ensemble layer depends on the predictions of the DAFT models, and requires solving either (i) a simple linear equation to optimize the weights of the ensemble layer, or (ii) training a $RF$ with shallow trees. Thus, only a single inference on each of the DAFT models with the few-shot samples is necessary  [^8] In our evaluations, we found that DAFT-E puts much more weight on the top few DAFT models. Hence if we have some training data from the target domain and have some budget constraints (e.g., the number of models), one solution could be to pick the models with the highest weights (from DAFT-E LR). Moreover, other transferability metrics can be used to select the top models within a budget. Some of those transferability metrics are LEEP score, MS-LEEP, E-LEEP, and SoftIoU-EEP (Agostinelli et al., 2022). LEEP scores can be calculated for each model individually (as discussed in Section 3). The other metrics, i.e., MS-LEEP, E-LEEP, and SoftIoU-EEP, are computed for a set of models and can be an alternative method to LR in DAFT-E to ensure better transferability for an ensemble of models.. In DA(FT)2, where we fine-tune a DAFT model, we need to perform computationally expensive back-propagation through the large model to update weights.
@@
[linecolor=black!0,backgroundcolor=orange!10]For few shot tasks, (cheap) weighted ensemble of DAFTs (DAFT-E) can outperform (expensive) DA(FT)2.
@@ 4.3 Overall Comparison
Figure 8: Performance comparison of DAFT-E ${}^{\text{\!\scriptsize Z}}$, DAFT-E, single-best DAFT and FFT. Error interval for DAFT-E is based on the random choice of the few-shot samples used to learn the weights of the ensemble aggregated over 10 trials. The single best DAFT for different datasets: Amazon (Roberta - Yelp), Cornell (Roberta - SST2), IMDB (xlnet - amazon), SST2 (xlnet - Cornell), Tweet (xlnet - SST2), Yelp (xlnet - amazon).
(a) MRPC
(b) QQP
(c) STSB
Figure 9: Performance comparison of DAFT-E ${}^{\text{\!\scriptsize Z}}$, DAFT-E, single-best DAFT and FFT for Text Similarity task. The error interval for DAFT-E is based on the random choice of the few-shot samples used to learn the weights of the ensemble, using RF method, aggregated over 10 trials. The single best DAFT for different datasets are: MRPC (Roberta - STSB), QQP (Roberta - STSB), STSB (Roberta - MRPC).
Figure 10: Performance of DAFT-E ${}^{\text{\!\scriptsize Z}}$, DAFT-E and Single Best DAFT for Medical Knowledge task with MMLU datasets.
For benchmarking purposes, let us assume that we have the entire training data ($D_{T}$) of the target domain, and have a model that is fine-tuned on $D_{T}$ (FFT on $D_{T}$). Let us compare the performance of the ensemble methods with this FFT on different datasets. Fig. 8 and 9 show the performance comparison of DAFT-E ${}^{\text{\!\scriptsize Z}}$ (green dash-dot line), DAFT-E (purple dash-dot line), single best DAFT (black dotted line), and FFT on $D_{T}$ (red dotted line) for all six datasets of the sentiment analysis and for all three datasets of the text similarity tasks respectively. From Fig. 8 and 9, we observe the advantage of using DAFT-E over DAFT-E ${}^{\text{\!\scriptsize Z}}$ when training data is available. It is interesting to note that DAFT-E has a strong increasing trend in performance (w.r.t. fine-tuning data size) for most cases, and catches up with either the single best DAFT or FFT for the target domain with the increase of fine-tuning data. The performance of DAFT-E compared to the best solution, i.e., FFT on the target domain, can be theoretically bound and is given by Proposition 2 in Section 5.
Table 4: Computational cost of methods shown in Fig. 1. $n$ denotes the number of few-shot samples. $C_{F}$ and $C_{B}$ denote the computational costs of a forward and backward pass for a LLM. $E$ is the number of fine-tuning epochs. $N$ is the number of DAFT models available to ensemble, $\bar{N}\leq N$ is the number of nonzero weights in the weighted ensemble.
@@
DAFT ${}^{\text{\!\scriptsize Z}}$ 00 $C_{F}$
DAFT-E ${}^{\text{\!\scriptsize Z}}$ 00 $N\cdot C_{F}$
@@
DAFT-E $Nn(C_{F}+E)$ $\bar{N}\cdot C_{F}$
Table 4 shows the overall comparison of the computational cost of all the five options that we have discussed in this paper. It should be noted that none of the ensemble methods suffer from the high computational complexity of back-propagation computation; further, using sparse weighting we can reduce the inference cost in DAFT-E compared to DAFT-E ${}^{\text{\!\scriptsize Z}}$. It is straightforward to see that DAFT ${}^{\text{\!\scriptsize Z}}$ and DAFT-E ${}^{\text{\!\scriptsize Z}}$ are the only ones among the five options discussed here that are zero-shot, and hence incurs zero training cost. In contrast, FT and DA(FT)2 has fine-tuning cost, i.e., back-propagation cost ($C_{B}$), along with forward propagation cost ($C_{F}$). For DAFT-E, the training cost is the forward propagation cost of $N$ DAFT models along with the linear cost for learning the weight of the ensemble layer. Lastly, FT, DAFT ${}^{\text{\!\scriptsize Z}}$ and DA(FT)2 use only a single (DAFT) model, and hence incurs only a forward propagation cost ($C_{F}$), whereas DAFT-E ${}^{\text{\!\scriptsize Z}}$ and DAFT-E use $N$ and $\tilde{N}$ DAFT models respectively in the inference stage.
To conclude the overall comparison, we present the performance of DAFT-E ${}^{\text{\!\scriptsize Z}}$, DAFT-E, and the single best DAFT model for the Medical Knowledge task using the MMLU datasets in Fig. 10. For DAFT-E, each dataset was split in half: one half was used to tune the linear weights of DAFT-E, and the other half was used for performance evaluation [^9] DAFT-E performance reported here uses the same set of hyper-parameters for ensemble-weight learning across all datasets; dataset-specific tuning could further improve results.. Since DAFT-E ${}^{\text{\!\scriptsize Z}}$ and DAFT models required no tuning, only the second half of the data was used for evaluation. This ensured that all three methods were evaluated on the same data. Due to the small size of each dataset, we repeated the random split 200 times for each of the 9 datasets and report the average performance. From the figure, we observe that DAFT-E outperformed DAFT-E ${}^{\text{\!\scriptsize Z}}$ in 7 out of the 9 tasks. Most notably, DAFT-E even outperformed the single best DAFT model in 5 of the 9 cases, demonstrating the effectiveness of the DAFT-E approach.
{mdframed}
[linecolor=black!0,backgroundcolor=orange!10]DAFT-E ${}^{\text{\!\scriptsize Z}}$ and DAFT-E are strong lightweight methods for leveraging already available DAFT models, often matching the single best DAFT model (only known post-hoc), and at times match the best possible performance of full fine-tuning with a large fine-tuning set.
4.4 Comparison with Model Soup
[…]
Dataset Uniform Soup DAFT-E ${}^{\text{\!\scriptsize Z}}$
Amazon 93.29% 93.72%
Cornell 85.09% 85.37%
IMDB 92.09% 92.28%
SST2 88.07% 89.55%
Tweet 86.09% 87.33%
Yelp 94.67% 95.30%
Table 6: Robustness Comparison of DAFT-E ${}^{\text{\!\scriptsize Z}}$ and Uniform Soup in terms of the set of models being combined. The results show that as we increase the number of DAFT models, the performance of both Uniform Soup and DAFT-E ${}^{\text{\!\scriptsize Z}}$ improves, with DAFT-E ${}^{\text{\!\scriptsize Z}}$ consistently outperforming Uniform Soup in the zero-shot setting. As discussed in Table 6, we are able to add many more models into DAFT-E ${}^{\text{\!\scriptsize Z}}$ than in Uniform Soup because Uniform Soup requires all models in the ensemble to have matching architectures.
DAFT models combined Uniform Soup DAFT-E ${}^{\text{\!\scriptsize Z}}$
SST2, Tweet 88.78% 89.09%
SST2, Tweet, Yelp 90.57% 90.86%
SST2, Tweet, Yelp, Cornell 91.19% 91.39%
SST2, Tweet, Yelp, Cornell, Amazon 92.09% 92.28%